4. At December 31, 2020, the following balances existed for AAA Corporation: Bonds Payable (4%) $500,000 Discount on Bonds Payable 10,000 The bonds mature on 12/31/27. Straight-line amortization is used. If 30% of the bonds are retired at 105 on January 1, 2024, what is the gain or loss on early extinguishment? Answer $_______________ Required: Compute the answer for each of the four problems. Show supporting computation. No need to show questions.
Answer #1
Cell Reference Particulars Amount $
A Face value of bond outstanding as at December 31,2020       500,000
B % of bonds early retired 30%
C=A*B Value of bond retired       150,000
D=C*105% Bond being retired at an value of $ 105       157,500
E=C-D Loss on early retirement of bonds         (7,500)
F 30% of discount on bond payable           3,000
G=E-F Loss on early retirement of bonds        (4,500)